OpenStudy (anonymous):

HOW TO SOLVE DIVISION PROBLEMS THAT HAVE EXPONENTS LIKE 11D^2/33D^4

OpenStudy (saifoo.khan):

\[\frac{11D^2}{33D^4} = \frac{1D^2}{3^4} = \frac{1}{3^{4-2}}\]

OpenStudy (saifoo.khan):

\[\frac{1}{3^2} = \frac{1}{9}\]

OpenStudy (anonymous):

The exponent is on D. So it is 1/3 * (1/D^2)

OpenStudy (saifoo.khan):

o_O